OpenStudy (anonymous):

Factor: x2 - 24x + 144

OpenStudy (maheshmeghwal9):

\[x^2-24x+144=x^2-12x-12x+144=(x-12)^2\]

OpenStudy (maheshmeghwal9):

whole sq. concept

OpenStudy (anonymous):

thank you! for verifying!

OpenStudy (maheshmeghwal9):

wc!